OpenStudy (anonymous):

(Linear Equations) Find the y-intercept for the equation: 3x + 2y = 9

OpenStudy (physicist):

Y-intercept is the y value where the graph crosses the y-axis, meaning that x=0. 3x +2y=9 3(0)+2y=9 2y=9 y=9/2 or 4.5